A 747 has a passenger cabin the is 60m x 15m. The plane is heading 20° N of E, at 450 km/h, while a wind blows 75 km/h Northwest. A bullet is fired from the front left corner of the plane stiking the backright corner in 0.20s. Find the Velocity of the bullet with respect to the earth
1
Expert's answer
2019-09-26T14:53:43-0400

The speed of bullet relative plane:


"v_{bp}=\frac{\sqrt{15^2+60^2}}{0.2}=309\frac{m}{s}"

The speed of bullet relative ground:

"v_{bg}=450-309=141\frac{m}{s}"